# Towards Privacy-Guaranteed Label Unlearning in Vertical Federated Learning: Few-Shot Forgetting without Disclosure

### v2

@@ Appendix A Appendix
This appendix provides additional theoretical insights, extended experimental results and further implementation details to support the main submission. We begin with a theoretical analysis of unlearning effectiveness in Sect. A.1, followed by a table of notations in Sect. A.2 for reference. Sect. A.3 provides an explanation of label leakage during the unlearning process in VFL, while Sect. A.4 discusses why our method can achieve strong results with only a small number of data samples. In addition, Sect. A.5 presents ablation experiments on applying our method to privacy-preserved VFL models. Sect. A.6 presents the experimental results for two-label and multi-label unlearning scenarios, showcasing the effectiveness of our solution in handling more complex unlearning tasks. This section also includes a detailed performance analysis of each baseline across three key metrics: (i) accuracy on $\mathcal{D}^{r}$, (ii) accuracy on $y^{u}$ and (iii) ASR score. Sect. A.7 presents a comparative table of our method against existing Federated Unlearning approaches, highlighting the capability of our method to preserve label privacy during the unlearning process. In Sect. A.8, we provide detailed information about the experiment setup, including model architectures, unlearning scenarios, baseline methods, datasets, and the data distribution among parties. Finally, in Sect. A.9, we discuss the limitations of our work, including aspects beyond its scope, methodological constraints, and other relevant considerations.
@@ A.2 Table of Notation
Notation Meaning $F_{\omega},G_{\theta_{k}}$ Active model and $k_{th}$ passive model $w^{u},\theta^{u}_{k}$ Unlearned active model and unlearned $k_{th}$ passive model $K$ The number of passive party $\lambda$ Mixed coefficient $\eta$ Learning rate $N$ Unlearning epochs $\mathbf{x}_{k}$ Private features own by $k_{th}$ passive party $y$ Private label owned by active party $y^{u}$ The unlearn labels $\{x_{k}^{u}\}$ The unlearned feature for client $k$ corresponding to the $y^{u}$ $x_{k}^{p}$ The known features for client $k$ corresponding to the $y^{u}$ $H_{k}$ Forward embedding of passive party $k$ $\vec{H}_{k}^{u},\vec{H}_{k}^{r}$ Augmented forward embedding of passive party $k$ of unlearn data and remain data. $g_{k}^{\prime}$ Gradient on the embedding $H_{k}^{\prime}$.
@@ A.4 Why Tiny Public Set Suffices
    Variance reduction via manifold mixup. Manifold mixup augments each public example into an infinite family of virtual points, reducing the gradient-estimator variance proportionally to the number of synthetic mixtures. Recent few-shot studies show that manifold-mixup greatly reduces sample requirements. For example, (Mangla et al., 2020) achieves a higher accuracy than prior baselines on four vision benchmarks using only 5 to 20 samples per class. Similarly, SimpliMix (Yang et al., 2024) achieves strong gains in few-shot 3-D point-cloud classification.
@@
    Tiny update magnitude. The pre-unlearning model already fits the data (loss $\approx 0$). As shown by (Chen et al., 2019), when the source and target domains are similar, even simple few-shot methods can perform well with minimal fine-tuning. Hence, once the gradient direction is reliable, only a small update is needed. In other words, because unlearning starts from a model that already fits the data, only the label-specific contribution must be removed, making the required parameter displacement is tiny. In the quadratic loss setting analysed in our theoretical analysis (Appendix A.1), the required step length scales with the gradient norm. Once the direction is correct, even a single mini-batch update can eliminate the label.
@@ A.5 Evaluation for different privacy preserving VFL methods
We evaluate our unlearning method under two privacy preserving VFL methods: (i) Differential Privacy (Fu et al., 2022b) and (ii) Gradient Compression (Fu et al., 2022b). Fig. 6 presents the effectiveness of our solution on both methods across different levels of variance Gaussian noise and compression ratio. Higher Gaussian noise levels and greater gradient compression ratios enhance privacy protection in VFL but lead to increased performance degradation. It shows that even for a large compression ratio and noise level, our proposed method can still unlearn effectively, while the utility of the vertical training decreases significantly.
@@ A.6.1 Utility Guarantee
From Tab. 3, the FT method shows performance consistent with the single-label unlearning scenario. It preserves $\mathcal{D}^{r}$ well on datasets with a small number of labels (e.g., CIFAR-10) but performs poorly on datasets with many labels (e.g., CIFAR-100). The Fisher Forgetting method consistently demonstrates poor preservation of $\mathcal{D}^{r}$ across all datasets and model architectures. The Amnesiac approach performs similarly to FT, achieving strong preservation of $\mathcal{D}^{r}$ on datasets with fewer labels (e.g., CIFAR-10) but showing reduced performance on datasets with a larger label space (e.g., CIFAR-100). The Unsir and BU methods also follow this trend, maintaining good preservation on simpler datasets like CIFAR-10, but experiencing greater degradation in $\mathcal{D}^{r}$ accuracy across all datasets compared to FT and Amnesiac. SSD demonstrates strong preservation of $\mathcal{D}^{r}$ across all experiments, with the exception of the CIFAR100 dataset under the VGG16 architecture.
@@
In contrast, our method demonstrates strong unlearning utility and achieves near-perfect preservation of $\mathcal{D}^{r}$ across all experimental settings, even in multi-label unlearning scenarios and for datasets with a large number of labels.
In conclusion, our experimental results across single-label, two-label, and multi-label unlearning tasks demonstrate that our method consistently delivers the best balance between utility and privacy preservation. For example, unlike FT, which maintains $\mathcal{D}^{r}$ accuracy only on datasets with few labels (e.g., CIFAR-10) but suffers on CIFAR-100, our method achieves near-perfect preservation of $\mathcal{D}^{r}$ regardless of label complexity. Fisher Forgetting performs poorly across all datasets and architectures, leading to significant drops in retained accuracy. Amnesiac Unlearning introduces instability through random relabeling of $\mathcal{D}_{u}$, causing accuracy degradation in high-label datasets, while Unsir and BU show even greater declines, especially in multi-label settings. SSD shows good preservation of $\mathcal{D}^{r}$ across different settings but its performance is slightly inferior to ours. In contrast, our method maintains high $\mathcal{D}^{r}$ accuracy (often above 98%) while reducing unlearned data accuracy close to random chance, consistently across all scenarios. These results provide strong evidence that our approach offers superior robustness, effectiveness, and generalizability for practical, privacy-preserving unlearning.
@@ A.6.2 Unlearning Effectiveness
From Tab. 4, the FT and Fisher Forgetting methods exhibit similarly poor unlearning effectiveness in the multi-label unlearning scenario, comparable to their performance in the two-label unlearning context. These methods are effective only in the single-label unlearning scenario. In contrast, the Amnesiac and Unsir methods consistently achieve perfect unlearning effectiveness in the multi-label scenario, mirroring their performance in single-label and two-label contexts. The BU method performs similarly to the two-label scenario in the multi-label context, showing limited unlearning effectiveness. SSD shows good unlearning effectiveness on multi-label scenario. Meanwhile, our method continues to demonstrate exceptional unlearning effectiveness, maintaining outstanding performance in multi-label scenarios, further reinforcing its robustness and efficiency.
@@
From Tab. 3 and 4, FT demonstrates excellent performance, achieving ASR scores that are consistently lower but close to those of the retrain model across all datasets, with the exception of the multi-labels scenario involving the ResNet18 architecture and the CIFAR100 dataset. In that case, the ASR score is only slightly higher than that of the retrain model. In contrast, Fisher Forgetting performs poorly in both the two-labels and multi-labels scenarios, exhibiting high ASR scores across all experiments, which indicates its ineffectiveness in mitigating adversarial risks. The Amnesiac method shows a consistent pattern of very low ASR scores across all experiments, potentially suggesting vulnerabilities to the Streisand effect, as observed in previous analyses. The Unsir method performs similarly to Fisher Forgetting, also achieving high ASR scores across all experiments, further highlighting its ineffectiveness. Boundary Unlearning continues to exhibit inconsistent performance throughout the experiments, failing to deliver reliable results. SSD continues to exhibit its good performance throughout the experiments, with consistently achieving ASR scores lower than and close to those of the retrain model. On the other hand, our method demonstrates outstanding performance on many datasets, achieving ASR scores that are consistently lower than and close to those of the retrain model. This highlights our method’s effectiveness and reliability in addressing adversarial risks while maintaining robust unlearning outcomes.
In conclusion, the results clearly demonstrate the effectiveness and robustness of our method in mitigating adversarial risks. Across various datasets, our approach consistently achieves ASR scores that are not only low but often even lower than those of the retrained models, indicating strong unlearning without compromising security. While FT and SSD perform reasonably well in most scenarios, other methods such as Fisher Forgetting, Unsir, and Boundary Unlearning exhibit limited effectiveness, with high ASR scores and inconsistent performance. Although the Amnesiac method achieves very low ASR scores, it raises concerns about potential vulnerabilities, such as the Streisand effect. In contrast, our method strikes the optimal balance between unlearning effectiveness and adversarial robustness, establishing it as the most reliable and secure solution among all evaluated approaches.
@@ A.6.3 Further Analysis of Baselines
Fisher Forgetting: This method struggles across all dimensions, exhibiting poor preservation of $\mathcal{D}^{r}$ irrespective of the dataset or architecture. It also fails to achieve unlearning effectiveness in two-label and multi-labels unlearning scenario, with high adversarial success rates (ASR) that compromise its ability to address adversarial risks. This consistent underperformance underscores the method’s lack of robustness in diverse scenarios.
Amnesiac Unlearning: While the Amnesiac method demonstrates strong unlearning effectiveness across various scenarios, its utility preservation declines significantly when applied to datasets with a large number of labels, such as CIFAR-100. This is primarily due to its approach of randomly assigning incorrect labels to $\mathcal{D}_{u}$, which introduces unpredictable shifts in decision boundaries and undermines reliability in more complex settings. Additionally, the method consistently yields very low ASR scores across all experiments. This implies that, although the low ASR scores indicate strong unlearning, they may also point to a susceptibility to the Streisand effect, as observed in prior analyses.
@@
SSD: SSD demonstrates strong performance across all three evaluation metrics, namely accuracy on $\mathcal{D}^{r}$, accuracy on $y^{u}$ and ASR score, although it remains marginally worse than our method in each case. Moreover, SSD incurs a higher computational cost, requiring longer runtime compared to our approach (see Fig. 2 in Sect. 5.2.3), which further limits its practical efficiency.
Overall, the experimental results highlight the key limitations of existing methods, such as FT, Fisher Forgetting, Amnesiac, Unsir, BU and SSD in striking a balance between utility preservation, unlearning effectiveness, and low ASR scores. While some methods may perform well in one aspect, they often fall short in others, especially when faced with complex or multi-label scenarios. In contrast, our method consistently delivers strong unlearning performance, near-perfect utility preservation, and robust ASR scores across all datasets and settings. This comprehensive and consistent performance firmly establishes our approach as the most reliable, effective, and scalable solution for tackling diverse and challenging unlearning tasks.
@@ A.7 Related Works
Method Type of FL Unlearning Types Unlearning Targets Protection for Unlearned Data FedEraser (Liu et al., 2021) Horizontal Exact Client FRU (Yuan et al., 2023) Approximate Client FedRecovery (Zhang et al., 2023a) Approximate Client VeriFI (Gao et al., 2024) Approximate Client HDUS (Ye et al., 2024) Approximate Client KNOT (Su & Li, 2023) Approximate Client FedRecover (Cao et al., 2023a) Approximate Client Knowledge Distillation (Wu et al., 2022) Approximate Client Discriminative Pruning (Wang et al., 2022) Approximate Class MoDe (Zhao et al., 2024) Approximate Class Rapid Retraining (Liu et al., 2022) Exact Sample QuickDrop (Dhasade et al., 2023) Approximate Sample FedAU (Gu et al., 2024b) Approximate Class, Sample & Client Ferrari (Gu et al., 2024a) Approximate Feature Fast Retraining (Wang et al., 2024) Vertical Exact Passive Party SecureCut (Li et al., 2024) Approximate Instance & Passive Party Constraint Imposing (Deng et al., 2023) Approximate Passive Party Backdoor Certification (Han et al., 2025) Approximate Passive Party Ours Approximate Label
@@ A.8 Experiment Setup
This section provides a detailed information on our experimental settings. Tables 9 and 10 summarize the hyper-parameters for our unlearning method. Table 11 summarizes the model name, VFL framework settings, datasets and unlearn labels involved in each unlearning scenario.
@@
Fisher Forgetting (Golatkar et al., 2020a): The dataset $\mathcal{D}^{r}$ is being divided among $K$ parties and the fisher information matrix (FIM) is being used to inject Gaussian noise to perturb the VFL baseline models towards exact unlearning. For every backward propagation, once the gradient is calculated, each party inject Gaussian noise into their respective VFL baseline model to perturb the baseline model toward exact unlearning.
@@
SSD (Foster et al., 2024b): The dataset is divided among $K$ parties and each parties modify their local model weights with the gradients of full data and $\mathcal{D}_{u}$.
@@
Table 9: Hyperparameters used for single-label unlearning in our proposed solution. In this single-label unlearning scenario, all datasets are processed using the same optimization method, stochastic gradient descent (SGD), with consistent parameters: a batch size of 32, a weight decay of 5e-4, and a momentum of 0.9. The recovery rate, which is the learning rate during the recovery phase, matches the unlearning rate for all datasets. Additionally, a minimal number of unlearning epochs (a maximum of 30) and unlearning data samples ($\mathcal{D}^{p,u}$, capped at 40) are utilized. Across all experiments, the number of recovery data samples ($\mathcal{D}^{p,r}$) used in the recovery phase is uniformly set to 3.
@@
Table 10: Hyperparameters used for two-label and multi-label unlearning in our proposed solution.In these unlearning scenarios, all datasets are processed using the same optimization method, stochastic gradient descent (SGD), with consistent parameters: a weight decay of 5e-4, and a momentum of 0.9. For the two-label unlearning scenario, a batch size of 32 is used, while a batch size of 64 is employed for the multi-label unlearning scenario. The recovery rate, which is the learning rate during the recovery phase, matches the unlearning rate for all datasets. Additionally, a minimal number of unlearning epochs (a maximum of 25) and unlearning data samples ($\mathcal{D}^{p,u}$, capped at 40) are utilized. Across all experiments, the number of recovery data samples ($\mathcal{D}^{p,r}$) used in the recovery phase is uniformly set to 3.
@@
Figures 7 and 8 illustrate how image features of the dataset are being split among the passive party. In this setup, we consider a scenario involving two passive parties. For all image datasets, except ModelNet, each image’s features are evenly split into two halves, as illustrated in Fig. 7. As an example in CIFAR10 (i.e., class = bird), each passive party is assigned one half of the image (bird) features. This means that neither party has access to the full set of image features for any given data point, ensuring that the data remains fragmented and partially hidden from each individual party. This approach is specifically designed to enhance privacy and security by ensuring that no single party has access to the full information of the original image. This partitioning allows collaborative computation or machine learning tasks to be performed while upholding data confidentiality.
For the ModelNet dataset, a different approach is used to distribute data between the two passive parties as illustrated in Fig. 8. Specifically, each 3D mesh model is rendered into two distinct 2D images, each capturing a unique view of the object from different perspectives. Then, these two generated 2D views are assigned to separate passive parties, with each party exclusively receiving one view. As a result, no single party has access to both perspectives of the 3D mesh model, preserving information separation. This strategy not only safeguards data privacy but also enables collaborative analysis or training, ensuring that each party’s contribution remains independent and incomplete.
@@
MNIST dataset contains images of handwritten digits. MNIST dataset comprises 60,000 training examples and 10,000 test examples. Each example is represented as a single-channel image with dimensions of $\times28$ pixels, categorized into one of 10 labels.

### v3

@@ Appendix A Appendix
This appendix provides additional theoretical insights, extended experimental results and further implementation details to support the main submission. We begin with a theoretical analysis of unlearning effectiveness in Section A.1, followed by a table of notations in Section A.2 for reference. Section A.3 provides an explanation of label leakage during the unlearning process in VFL, while Section A.4 discusses why our method can achieve strong results with only a small number of data samples. In addition, Section A.5 presents ablation experiments on applying our method to privacy-preserved VFL models. Section A.6 presents the experimental results for two-label and multi-label unlearning scenarios, showcasing the effectiveness of our solution in handling more complex unlearning tasks. This section also includes a detailed performance analysis of each baseline across three key metrics: (i) accuracy on $\mathcal{D}^{r}$, (ii) accuracy on $y^{u}$ and (iii) ASR score. Section A.7 presents a comparative table of our method against existing Federated Unlearning approaches, highlighting the capability of our method to preserve label privacy during the unlearning process. In Section A.8, we provide detailed information about the experiment setup, including model architectures, unlearning scenarios, baseline methods, datasets, and the data distribution among parties. Finally, in Section A.9, we discuss the limitations of our work, including aspects beyond its scope, methodological constraints, and other relevant considerations.
@@ A.2 Table of Notation
Notation Meaning $F_{\omega},G_{\theta_{k}}$ Active model and $k_{th}$ passive model $w^{u},\theta^{u}_{k}$ Unlearned active model and unlearned $k_{th}$ passive model $K$ The number of passive party $\lambda$ Mixed coefficient $\eta$ Learning rate $N$ Unlearning epochs $\mathbf{x}_{k}$ Private features own by $k_{th}$ passive party $y$ Private label owned by active party $y^{u}$ The unlearn labels $\{x_{k}^{u}\}$ The unlearned feature for client $k$ corresponding to the $y^{u}$ $x_{k}^{p}$ The known features for client $k$ corresponding to the $y^{u}$ $H_{k}$ Forward embedding of passive party $k$ $\vec{H}_{k}^{u},\vec{H}_{k}^{r}$ Augmented forward embedding of passive party $k$ of unlearn data and remain data.
@@ A.4 Why Tiny Public Set Suffices
    Variance reduction via manifold mixup. Manifold mixup augments each public example into an infinite family of virtual points, reducing the gradient-estimator variance proportionally to the number of synthetic mixtures. Recent few-shot studies show that manifold-mixup greatly reduces sample requirements. For example, (Mangla et al., 2020) achieves higher accuracy than prior baselines across four vision benchmarks with only 5 to 20 samples per class. Similarly, SimpliMix (Yang et al., 2024) achieves strong gains in few-shot 3D point cloud classification.
@@
    Tiny update magnitude. The pre-unlearning model already fits the data (loss $\approx 0$). As shown by (Chen et al., 2019), when the source and target domains are similar, even simple few-shot methods can perform well with minimal fine-tuning. Hence, once the gradient direction is reliable, only a small update is needed. In other words, because unlearning starts from a model that already fits the data, only the label-specific contribution needs to be removed, resulting in a small parameter displacement. In the quadratic loss setting analysed in our theoretical analysis (Appendix A.1), the required step length scales with the gradient norm. Once the direction is correct, even a single mini-batch update can eliminate the label.
@@ A.5 Evaluation for different privacy preserving VFL methods
We evaluate our unlearning method under two privacy preserving VFL methods: (i) Differential Privacy (Fu et al., 2022a) and (ii) Gradient Compression (Fu et al., 2022a). Fig. 6 presents the effectiveness of our solution on both methods across different levels of variance Gaussian noise and compression ratio. Higher Gaussian noise levels and greater gradient compression ratios enhance privacy protection in VFL but lead to increased performance degradation. It shows that even for a large compression ratio and noise level, our proposed method can still unlearn effectively, while the utility of the vertical training decreases significantly.
@@ A.6.1 Utility Guarantee
From Tab. 3, the FT method shows performance consistent with the single-label unlearning scenario. It preserves $\mathcal{D}^{r}$ well on datasets with a small number of labels (e.g., CIFAR-10) but performs poorly on datasets with many labels (e.g., CIFAR-100). The Fisher Forgetting method consistently demonstrates poor preservation of $\mathcal{D}^{r}$ across all datasets and model architectures. The Amnesiac approach performs similarly to FT, achieving strong preservation of $\mathcal{D}^{r}$ on datasets with fewer labels (e.g., CIFAR-10) but showing reduced performance on datasets with a larger label space (e.g., CIFAR-100). The Unsir and BU methods also follow this trend, maintaining good preservation on simpler datasets like CIFAR-10, but experiencing greater degradation in $\mathcal{D}^{r}$ accuracy across all datasets compared to FT and Amnesiac. SSD demonstrates strong preservation of $\mathcal{D}^{r}$ across all experiments, except for the CIFAR100 dataset under the VGG16 architecture.
@@
In contrast, our method demonstrates strong unlearning utility and achieves near-perfect preservation of $\mathcal{D}^{r}$ across all experimental settings, including multi-label unlearning scenarios and for datasets with a large number of labels.
In conclusion, our experimental results across single-label, two-label, and multi-label unlearning tasks demonstrate that our method consistently delivers the best balance between utility and privacy preservation. For example, unlike FT, which maintains $\mathcal{D}^{r}$ accuracy only on datasets with few labels (e.g., CIFAR-10) but suffers on CIFAR-100, our method achieves near-perfect preservation of $\mathcal{D}^{r}$ regardless of label complexity. Fisher Forgetting performs poorly across all datasets and architectures, leading to significant drops in retained accuracy. Amnesiac Unlearning introduces instability through random relabeling of $\mathcal{D}_{u}$, leading to accuracy degradation on high-label datasets, whereas Unsir and BU show even greater declines, especially in multi-label settings. SSD shows good preservation of $\mathcal{D}^{r}$ across different settings, but its performance is slightly inferior to ours. In contrast, our method maintains high $\mathcal{D}^{r}$ accuracy (often above 98%) while reducing unlearned data accuracy close to random chance, consistently across all scenarios. These results provide strong evidence that our approach offers superior robustness, effectiveness, and generalizability for practical, privacy-preserving unlearning.
@@ A.6.2 Unlearning Effectiveness
From Tab. 4, the FT and Fisher Forgetting methods exhibit similarly poor unlearning effectiveness in the multi-label unlearning scenario, comparable to their performance in the two-label unlearning context. These methods are effective only in the single-label unlearning scenario. In contrast, the Amnesiac and Unsir methods consistently achieve perfect unlearning effectiveness in the multi-label scenario, mirroring their performance in single-label and two-label contexts. The BU method performs similarly to the two-label scenario in the multi-label context, showing limited unlearning effectiveness. SSD shows good unlearning effectiveness in a multi-label scenario. Meanwhile, our method continues to demonstrate exceptional unlearning effectiveness, maintaining outstanding performance in multi-label scenarios, further reinforcing its robustness and efficiency.
@@
From Tab. 3 and 4, FT demonstrates excellent performance, achieving ASR scores that are consistently lower but close to those of the retrain model across all datasets, with the exception of the multi-labels scenario involving the ResNet18 architecture and the CIFAR100 dataset. In that case, the ASR score is only slightly higher than that of the retrain model. In contrast, Fisher Forgetting performs poorly in both the two-labels and multi-labels scenarios, exhibiting high ASR scores across all experiments, which indicates its ineffectiveness in mitigating adversarial risks. The Amnesiac method shows a consistent pattern of very low ASR scores across all experiments, potentially suggesting vulnerabilities to the Streisand effect, as observed in previous analyses. The Unsir method performs similarly to Fisher Forgetting, also achieving high ASR scores across all experiments, further highlighting its ineffectiveness. Boundary Unlearning continues to exhibit inconsistent performance throughout the experiments, failing to deliver reliable results. SSD continues to perform well across the experiments, consistently achieving ASR scores lower than and close to those of the retrain model. On the other hand, our method demonstrates outstanding performance across many datasets, achieving ASR scores that are consistently lower than, and close to, those of the retrain model. This highlights our method’s effectiveness and reliability in addressing adversarial risks while maintaining robust unlearning outcomes.
In conclusion, the results clearly demonstrate the effectiveness and robustness of our method in mitigating adversarial risks. Across datasets, our approach consistently achieves ASR scores that are not only low but also often lower than those of the retrained models, indicating strong unlearning without compromising security. While FT and SSD perform reasonably well in most scenarios, other methods such as Fisher Forgetting, Unsir, and Boundary Unlearning exhibit limited effectiveness, with high ASR scores and inconsistent performance. Although the Amnesiac method achieves very low ASR scores, it raises concerns about potential vulnerabilities, such as the Streisand effect. In contrast, our method strikes the optimal balance between unlearning effectiveness and adversarial robustness, making it the most reliable and secure solution among the evaluated approaches.
@@ A.6.3 Further Analysis of Baselines
Fisher Forgetting: This method struggles across all dimensions, exhibiting poor preservation of $\mathcal{D}^{r}$ irrespective of the dataset or architecture. It also fails to achieve unlearning effectiveness in two-label and multi-labels unlearning scenarios, with high adversarial success rates (ASR) that compromise its ability to address adversarial risks. This consistent underperformance underscores the method’s lack of robustness in diverse scenarios.
Amnesiac Unlearning: While the Amnesiac method demonstrates strong unlearning effectiveness across various scenarios, its utility preservation declines significantly when applied to datasets with a large number of labels, such as CIFAR-100. This is primarily due to its approach of randomly assigning incorrect labels to $\mathcal{D}_{u}$, which introduces unpredictable shifts in decision boundaries and undermines reliability in more complex settings. Additionally, the method consistently yields very low ASR scores across all experiments. This implies that although low ASR scores indicate strong unlearning, they may also indicate susceptibility to the Streisand effect, as observed in prior analyses.
@@
SSD: SSD demonstrates strong performance across all three evaluation metrics, namely accuracy on $\mathcal{D}^{r}$, accuracy on $y^{u}$ and ASR score, although it remains marginally worse than our method in each case. Moreover, SSD incurs a higher computational cost, resulting in longer runtime than our approach (see Fig. 2 in Sect. 5.2.3), further limiting its practical efficiency.
Overall, the experimental results highlight the key limitations of existing methods, such as FT, Fisher Forgetting, Amnesiac, Unsir, BU and SSD in striking a balance between utility preservation, unlearning effectiveness, and low ASR scores. While some methods may perform well in one aspect, they often fall short in others, especially in complex or multi-label scenarios. In contrast, our method consistently delivers strong unlearning performance, near-perfect utility preservation, and robust ASR scores across all datasets and settings. This comprehensive and consistent performance firmly establishes our approach as the most reliable, effective, and scalable solution for tackling diverse and challenging unlearning tasks.
@@ A.7 Related Works
Method Type of FL Unlearning Types Unlearning Targets Protection for Unlearned Data FedEraser (Liu et al., 2021) Horizontal Exact Client FRU (Yuan et al., 2023) Approximate Client FedRecovery (Zhang et al., 2023a) Approximate Client VeriFI (Gao et al., 2024) Approximate Client HDUS (Ye et al., 2024) Approximate Client KNOT (Su and Li, 2023) Approximate Client FedRecover (Cao et al., 2023a) Approximate Client Knowledge Distillation (Wu et al., 2022) Approximate Client Discriminative Pruning (Wang et al., 2022) Approximate Class MoDe (Zhao et al., 2024) Approximate Class Rapid Retraining (Liu et al., 2022) Exact Sample QuickDrop (Dhasade et al., 2023) Approximate Sample FedAU (Gu et al., 2024b) Approximate Class, Sample & Client Ferrari (Gu et al., 2024a) Approximate Feature Fast Retraining (Wang et al., 2024) Vertical Exact Passive Party SecureCut (Li et al., 2024) Approximate Instance & Passive Party Constraint Imposing (Deng et al., 2023) Approximate Passive Party Backdoor Certification (Han et al., 2025) Approximate Passive Party Ours Approximate Label
@@ A.8 Experiment Setup
This section provides detailed information on our experimental settings. Tables 9 and 10 summarize the hyper-parameters for our unlearning method. Table 11 summarizes the model name, VFL framework settings, datasets and unlearn labels involved in each unlearning scenario.
@@
Fisher Forgetting (Golatkar et al., 2020a): The dataset $\mathcal{D}^{r}$ is being divided among $K$ parties, and the Fisher information matrix (FIM) is being used to inject Gaussian noise to perturb the VFL baseline models towards exact unlearning. For every backward propagation, once the gradient is calculated, each party inject Gaussian noise into their respective VFL baseline model to perturb the baseline model toward exact unlearning.
@@
SSD (Foster et al., 2024b): The dataset is divided among $K$ parties, and each party modifies their local model weights with the gradients of full data and $\mathcal{D}_{u}$.
@@
Table 9: Hyperparameters used for single-label unlearning in our proposed solution. In this single-label unlearning scenario, all datasets are processed using the same optimization method, stochastic gradient descent (SGD), with consistent parameters: a batch size of 32, a weight decay of 5e-4, and a momentum of 0.9. The recovery rate, which is the learning rate during the recovery phase, matches the unlearning rate across all datasets. Additionally, a minimal number of unlearning epochs (a maximum of 30) and unlearning data samples ($\mathcal{D}^{p,u}$, capped at 40) are utilized. Across all experiments, the number of recovery data samples ($\mathcal{D}^{p,r}$) used in the recovery phase is uniformly set to 3.
@@
Table 10: Hyperparameters used for two-label and multi-label unlearning in our proposed solution.In these unlearning scenarios, all datasets are processed using the same optimization method, stochastic gradient descent (SGD), with consistent parameters: a weight decay of 5e-4, and a momentum of 0.9. For the two-label unlearning scenario, a batch size of 32 is used, while a batch size of 64 is employed for the multi-label unlearning scenario. The recovery rate, which is the learning rate during the recovery phase, matches the unlearning rate across all datasets. Additionally, a minimal number of unlearning epochs (a maximum of 25) and unlearning data samples ($\mathcal{D}^{p,u}$, capped at 40) are utilized. Across all experiments, the number of recovery data samples ($\mathcal{D}^{p,r}$) used in the recovery phase is uniformly set to 3.
@@
Figures 7 and 8 illustrate how image features of the dataset are being split among the passive party. In this setup, we consider a scenario involving two passive parties. For all image datasets, except ModelNet, each image’s features are evenly split into two halves, as illustrated in Fig. 7. As an example in CIFAR10 (i.e., class = bird), each passive party is assigned one half of the image (bird) features. This means that neither party has access to the full set of image features for any given data point, ensuring that the data remains fragmented and partially hidden from each individual party. This approach is specifically designed to enhance privacy and security by ensuring that no single party has access to the full information of the original image. This partitioning enables collaborative computation or machine learning tasks while maintaining data confidentiality.
For the ModelNet dataset, a different approach is used to distribute data between the two passive parties as illustrated in Fig. 8. Specifically, each 3D mesh model is rendered into two distinct 2D images, each capturing a unique view of the object from different perspectives. Then, these two generated 2D views are assigned to separate passive parties, with each party receiving one view exclusively. As a result, no single party has access to both perspectives of the 3D mesh model, preserving information separation. This strategy not only safeguards data privacy but also enables collaborative analysis or training, ensuring that each party’s contribution remains independent and incomplete.
@@
MNIST is a dataset of handwritten digits. MNIST dataset comprises 60,000 training examples and 10,000 test examples. Each example is represented as a single-channel image with dimensions of $\times28$ pixels, categorized into one of 10 labels.